Use what you need of the following set of information to answer the question:

S0 = $1.1420/€            S0 = £0.7776/$                        S0 = $1.2860/£       

                        E (S1 (3-months)) = $1.1280/€                              F3-months = $1.1590/€

                 Where: € = euro; £ = British pound; E ( ) = Expected.

4.    If one year ago the spot price of British pounds in terms of the $ was $1.3680/£. Do you think the dollar has appreciated or depreciated against the British pound over the last year (explain answer without any calculations)? What was the percentage rate of change in the value of the dollar over the last year? Is your numerical answer consistent with your explanation?

(Show all your calculations and explain your work. All currency amount answers should be presented with two decimal places, all currency rates with the same number of decimals they are given, and all rates of change and premiums/discounts with four decimal places)
Answer #1

The spot price of British Pound in terms of US Dollar was 1.3680 US Dollar per GBP.

Currently the price of British Pound in terms of US Dollar is 1.2860 US Dollar per GBP.

So the number of US Dollars per GBP has reduced. That means the US Dollar has appreciated in value against the British pound.

1 Dollar ( 1 year ago) = 1/1.3680 = 0.73099 GBP

1 Dollar (Today) = 0.7776 GBP

So change in value of Dollar = (0.7776 - 0.73099)/(0.73099) = 0.06376


% change = 6.376%